OpenStudy (anonymous):

Find dw/dt without converting w to a function of t and using the appropriate chain rule W = xycosz; x = t; y = t^2 ; z= arccost I don't think this is possible

OpenStudy (inkyvoyd):

convert W into a function of t. Any ideas?

OpenStudy (anonymous):

dw/dt = dw/dx dx/dt * dw/dy dy/dt * dw/dz dz/dt ?

OpenStudy (anonymous):

But there is an arccos 1/sqrt(). ohh nuuuu D:

OpenStudy (inkyvoyd):

my bad, I can't read for shiz.

OpenStudy (inkyvoyd):

Is this a multivariate calculus quesiton?

OpenStudy (anonymous):

yes :(

OpenStudy (inkyvoyd):

bahaha, I'm not qualified to do this, but I'll give what I'd do: dW/dt= dw/dx dx/dt+dw/dy dy/dt+dw/dz dz/dt

OpenStudy (inkyvoyd):

as for the arccosine you automatically assume that domain of z is within that of the arccosine , so the W(x,y,z) cancels out for jus the partial z

OpenStudy (anonymous):

oh. Thanks for the advice

OpenStudy (inkyvoyd):

dz/dt is a formula that you should have memorized. d/dx arccos(x)=-1/sqrt(1-x^2)

OpenStudy (anonymous):

That's the what the teacher said too

OpenStudy (inkyvoyd):

what's the main issue you are seeing?

OpenStudy (anonymous):

partial derivitives scare me, I never do them right ;/

OpenStudy (anonymous):

I get an angle that i put into the arccos and it's crazy bad. But the teacher said they would cancle too. So you guys must know what's up.

OpenStudy (anonymous):

i've got the answer It's hard :/

OpenStudy (inkyvoyd):

well, just follow the rule. dW/dt= dw/dx dx/dt+dw/dy dy/dt+dw/dz dz/dt assuming you already have the first two terms of that down, we'll just look at dw/dz dz/dt dw/dz=-xy sin z; dz/dt=-1/sqrt(1-t^2) dw/dz dz/dt=\(\Huge \frac{-xy\sin z}{\sqrt{1-t^2}}\) we know that x=t y=t^2 z=arccos(t) so dw/dz dz/dt=\(\Huge \frac{-t^3\sin z}{\sqrt{1-t^2}}\) at this point recall that sin^2 theta+cos^2 theta=1, but also that z must be between 0 and pi inclusive, so sin z must be between 0 and 1 (sin z is positive). Therefore we shall take the positive square root of sin z=sqrt(1-cos^2 t)

OpenStudy (inkyvoyd):

*cos^2 z

OpenStudy (anonymous):

Thank you for your help though if your curious it's w = t^4, dw/dt = 4t^3

OpenStudy (anonymous):

oh, that's the same thing?

OpenStudy (inkyvoyd):

\(\Huge \frac{-t^3\sqrt{1-\cos^2(\arccos t)}}{\sqrt{1-t^2}}=-t^3\)

OpenStudy (inkyvoyd):

what's the same thing?

OpenStudy (anonymous):

is that equivilent to the answer ;/ arccos is very hard for me to understand.

OpenStudy (anonymous):

but it should be 4t^3

OpenStudy (inkyvoyd):

Well, I only evaluated dw/dz dz/dt

OpenStudy (inkyvoyd):

I believe I got the sign wrong, and should've gotten a t^3

OpenStudy (inkyvoyd):

dw/dz=-xy sin z; dz/dt=-1/sqrt(1-t^2) yes, in this line they should cancel to make positive.

OpenStudy (inkyvoyd):

If you want, I am perfectly willing to start over and go through the entire problem with you the way it was supposed to be done without doing the substitutions for t

OpenStudy (anonymous):

;D omg that would be great. But I must warn you i'm horrible at math

OpenStudy (inkyvoyd):

okay. So let's start over; I'll recite the chain rule. \(\Huge \frac{dW}{dt}=\frac{\partial W}{\partial x} \frac{dx}{dt}+\frac{\partial W}{\partial y}\frac{dy}{dt}+\frac{\partial W}{\partial z}\frac{dz}{dt} \)

OpenStudy (anonymous):

i reamber that because it always criss crosses with the x, then y, then the z :( but so hard to use

OpenStudy (anonymous):

and I think arcsin and arccos are always used to find theda

OpenStudy (inkyvoyd):

As we can see it's a three part problem, so let's just start with the x-portion, then the y-portion, then the z-portion. Calculating dw/dx we get simply y cos z because y cos z is a constant, and x becomes 1 From the problem, y=t^2, and z=arccos t Therefore dw/dx=(t^2)cos(arccos t)=t^3 Calculating dx/dt we get 1 therefore the first term \(\Huge\frac{\partial W}{\partial x} \frac{dx}{dt}=t^3*1=t^3\)

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

Now for the second part. Calculating dw/dy we get simply x cos z because x cos z is a constant, and y becomes 1. From the problem, x=t, and z=arccos t Therefore dw/dy=(t)cos(arccos t)=t^2 Calculating dy/dt we get 2t. Therefore the second term \(\Huge \frac{dW}{dt}=\frac{\partial W}{\partial y}\frac{dy}{dt}=t^2*2t=2t^3\)

OpenStudy (anonymous):

adding them together would give us 5t^3?

OpenStudy (inkyvoyd):

Now for the third part, the trickiest part. Calculating dw/dz we get -xy sin(z) because xy is a constant, and cos(z) becomes - sin z. From the problem, x=t, y=t^2, therefore dw/dz=-t^3 sin(z) Notice I have not gotten rid of the z. since it's not immediately apparent what to do with it, it's best to leave it alone for now. Calculating dz/dt we get -1/sqrt(1-t^2) The trick here is to recall the pythagorean identity from trigonometry, that is \(\Large \sin^2\theta+\cos^2\theta=1\) Because of this, \(\Large \sin\theta=\pm \sqrt{1-\cos^2\theta}\) (looks so similar to dz/dt) But this is an identity, and thus applies to any variable; in this case z is of interest because we got dw/dz=-t^3 sin(z) earlier; \(\Large \sin z=\pm \sqrt{1-\cos^2 z}\) The big question is, do we use the plus or the minus? The answer lies within analyzing the domain and range of the functions we were originally given.

OpenStudy (anonymous):

why do we look in the domain? because it has to be a real number?

OpenStudy (inkyvoyd):

recall that z=arccos(t), and that W=xy cos(z) The range of the arccos function is between 0 and pi. Remember that we are interested in taking the sin of z. In other words, we want to look at all possible values of the sin of z to determine whether or not the sign is positive (if we know the sin of z is always positive, then we take the positive square root because the square root yields a positive number; if we know that the sin of z is negative, we take the negative square root for hte same reason). However, we know that the z=arccos(t), meaning that we are taking the sin of z from 0 to pi. If you look at a graph of sin z, you'll find that sin z from [0,pi] is really the first "bump" of the graph. Thus because sin z is always positive, we know that our expression \(\pm\sqrt{1-\cos^2 z}\) is always positive, and because the square root of a number must be positive (in the real range), we must take the positive root because plus times plus makes plus. Thus we know that \(\Huge \sin z=+\sqrt{1-\cos^2 z}\)